# Also at ] The European Space Agency Centre for Earth Observation, Frascati, Italy

### v3

@@ VI.2 Appendix B: Derivation of advective transport
In this section, we explicitly derive the advective transport term for the model, providing closure to the argument that active chiral matter can exhibit Euler turbulence.
@@
To close the hydrodynamic equations, we compare the microscopic and macroscopic tensors. In order to do so, we posit a constitutive relation of the form $\langle v_{i}v_{j}\rangle=P_{\text{eff}}\sigma_{ij}+\lambda u_{i}u_{j}$, where $\lambda$ scales the advective term. Substituting the expansions from Eqs. (23) and (25) into this ansatz for the $xx$-component yields:
@@
Consequently, our effective pressure $P_{\text{eff}}\propto(1-R^{4})$ decays much more slowly with ordering than the standard $(1-R^{2})$ prediction. In our model, this creates a thermodynamic barrier: the isotropic ’disorder pressure’ remains dominant even as the system attempts to synchronize. This explains the stability of the active vortex glass (Phase II): the intermediate agitation ($\Delta\omega$) is insufficient to overcome this pressure plateau, causing the system to jam in a disordered state ($R\to 0$). Only the high agitation of Phase III provides the energy density required to penetrate this barrier and sustain the regime of marginal synchronization ($0<R<1$) characteristic of the Onsager condensate.
@@
where we identify $F_{active}\approx\langle\sum\dot{v}_{i}\delta\rangle$ from Eq. (19) as a force vector that arise from the intrinsic frustration ($\dot{\mathbf{v}}_{i}$ depends on angular velocity $\dot{\phi}_{i}\approx\omega_{i}$).
The shallow water analogy
@@
As stated in the Methodology section, the interpretation of Eq. (12) identifies the system being modeled as formally analogous to the shallow water framework, where the active density $\rho$ acts as the fluid height. However, we must point out a nuance introduced by a critical thermodynamic distinction in the equation of state. In classical shallow water systems, hydrostatic pressure scales with height ($P\propto\rho^{2}$), leading to dispersion [23]. In Eq. (29), we identified the pressure term $P_{\text{eff}}\propto(1-R^{4})$, meaning that the effective pressure scales with disorder. Crucially, because the local order parameter $R$ arises from the mean-field convolution of the particle density (Eq. 1), regions of high density support high synchronisation ($R\sim\rho$; realized by $R$ being the magnitude of the complex mean field, in turn a measurement of $\rho$, since phasors align in local synchronisation regions). Consequently, the pressure gradient $\nabla P_{\text{eff}}\propto-\nabla R^{4}$ acts as a suction force, drawing agents from disordered regions ($R\approx 0$) toward ordered vortex cores ($R\approx 1$). This manifests as a fluid with effective negative compressibility, driving the system to phase-separate into the coherent topological structures observed in the Onsager dipole, rather than dispersing into noise.
@@
Consequently, the inviscid Euler behavior that we observe in the RFE field represents the low-Mach-number limit of this compressible system, where, we reiterate, the rotational, non-compressive wave modes dominate the irrotational, compressive wave (acoustic) modes at large scales. Hence, in the hydrodynamic limit (the Onsager dipole, or Phase III in Table 1), we recover in Eq. (30) the effective Euler advection term $\lambda(\mathbf{u}\cdot\nabla)\mathbf{u}$.
@@
It is instructive to contrast the foregoing with the defect-driven accumulation observed in active nematics [49] and flow-stabilized vortex lattices [50]. In those systems, mass transport is driven by the divergence of an active stress that scales linearly with the nematic tensor order parameter ($\sigma\sim\mathbf{Q}$), which creates effective pressure wells that scale akin to $\propto(1-R^{2})$.
In our model, however, the emergent inertia of the topological gas ($\lambda=R^{2}$) yields a stiffer effective pressure $P_{eff}\propto(1-R^{4})$ (Eq. 29). This creates sharper, ’hard-core’ density cavities at defect sites (visible in Figure 3), providing additional justification for the appropriateness of the point vortex (Kirchhoff-Onsager) Hamiltonian in topological gas dynamics, since it assumes distinct, non-overlapping cores [19].
The role of dispersion
Finally, we clarify the role of the dispersion $\Delta\omega$. The magnitude of the active Reynolds stress scales with the velocity variance $\langle\delta v^{2}\rangle$. Since the velocity fluctuations in the fluid limit are driven by the deviation of $\phi_{i}$ from $\Phi$,
Finally, we clarify the role of the dispersion $\Delta\omega$. The magnitude of the active Reynolds stress $\tau_{\text{active}}$ scales with the velocity variance $\langle\delta v^{2}\rangle$. To see why, observe that fluctuation in velocity $\delta v_{i}$ is driven by the deviation of the particle’s natural frequency $\omega_{i}$ from the local mean frequency $\Phi$, $\delta v_{i}\sim\delta\dot{\phi}_{i}\approx(\omega_{i}-\Phi)$, meaning that,
$\langle\delta v\delta v\rangle\propto\langle(\omega_{i}-\Phi)^{2}\rangle.$ (31)
By definition, the variance of the natural phases $\omega_{i}$ is the dispersion parameter squared:
$\tau_{\text{active}}\propto\langle(\omega_{i}-\Phi)^{2}\rangle\approx(\Delta\omega)^{2}.$ (32)
Thus, $\Delta\omega$ functions as the agitation parameter or granular temperature of the topological gas. This creates the central thermodynamic trade-off of the system:
(1) Low $\Delta\omega$ (Phase I): High synchronisation ($R\to 1$) but zero stress ($\tau_{active}\to 0$). The fluid freezes into static clumps.
(2) High $\Delta\omega$ (Phase III): High stress ($\tau_{active}$ large) and marginal synchronisation ($0<R<1$). The system generates sufficient active stress to drive the inverse cascade, creating the “phantom inertia” that sustains the Onsager dipole.
Conversely, in Phase II, the active vortex glass, agitation $\Delta\omega$ is insufficient to allow the system to overcome the lattice defect and access the negative-temperature regime of the Onsager dipole. Thus, ’heating’ the microscopic degrees of freedom paradoxically leads to macroscopic ordering, a signature of non-equilibrium phase transitions where noise-induced coherence allows the emergence of large-scale transport [18].

### v4

@@ VII Supplementary Materials: Derivation of the hydrodynamic limit
In this section, we derive the hydrodynamic limit of our model, beginning with deriving the macroscopic velocity, the advective transport term for the model, and establishing an analogy with shalow water hydrodynamics, providing closure to the argument that active chiral matter can exhibit Euler turbulence. We note that we reinstate the units in this section to track dimensional consistency.
VII.1 Velocity and continuity
We define the macroscopic velocity field $\mathbf{u}(\mathbf{r},t)$ as the local average of the particle velocities. Decomposing the particle phase into the local mean phase $\Psi$ and a fluctuation $\delta\phi_{i}$ (that is, $\phi_{i}=\Psi+\delta\phi_{i}$), and assuming symmetric fluctuations (that is, $\langle\sin\delta\phi\rangle\approx 0$), the coarse-grained velocity becomes,
$\mathbf{u}=\langle\mathbf{v}_{i}\rangle\approx v_{0}\langle\cos\delta\phi\rangle(\cos\Psi,\sin\Psi).$ (12)
We identify the local order parameter $R_{1}$,
for a peaked (narrow) Gaussian distribution in $\delta\phi$; we then drop the subscript. $R$ measures the degree of of local synchronisation. Thus, we obstain,
$|\mathbf{u}|=v_{0}R.$ (14)
While individual particles move at constant speed $v_{0}$, the renormalised fluid element speed is variable and proportional to the local synchronisation. To express $\mathbf{u}$ in terms of the mean field $\Psi$, we identify the flow direction $\mathbf{\hat{n}}=(\cos\Psi,\sin\Psi)$ with the normalized phase gradient:
$\mathbf{u}=v_{0}R\frac{\nabla\Psi}{|\nabla\Psi|}.$ (15)
$\mathbf{u}\approx R\nabla\Psi.$ (16)
This approximation eventually allows us to utilize the machinery of inviscid fluid mechanics in the bulk of the domain, while the factor $R$ naturally regularizes the dynamics at the topological singularities.
Before providing closure to the above, we need to address continuity. The continuity equation for $\mathbf{u}$ then reads,
$\frac{\partial\rho}{\partial t}+\nabla\cdot(\rho\mathbf{u})=0.$ (17)
This formulation identifies the system as formally analogous to the shallow water equations [65, 66], where the active density $\rho$ acts as the fluid height (see Appendix-subsection VII.3). The phase curvature $\nabla^{2}\Psi$ drives the compression of the fluid elements, explicitly linking topological defects to density fluctuations.
Eq. (8) implies $|\nabla\Psi|=m/r$. Substituting this into our macroscopic velocity definition (Eq. 15), and assuming the order parameter $R$ is locally uniform, the fluid velocity induced by a defect scales as:
$|\mathbf{u}|\approx v_{0}R\frac{m}{r}.$ (19)
This confirms that the topological defects of the mean phase act as point vortices with an effective physical circulation $\Gamma\approx 2\pi mv_{0}R$. Consequently, we arrive at the Kirchhoff-Onsager Hamiltonian in Eq. (9). Crucially, this derivation highlights that the topological charge driving the Hamiltonian dynamics is scaled by the order parameter $R$. If renormalization (or synchrony) is suppressed ($R\to 0$), the effective circulation vanishes ($\Gamma_{\text{eff}}\to 0$), causing the Hamiltonian interaction to collapse.
VII.2 Momentum and Pressure
@@
To relate the foregoing to the hydrodynamic equations, we compare the microscopic and macroscopic tensors. In order to do so, we posit a constitutive relation of the form $\langle v_{i}v_{j}\rangle=P_{\text{eff}}\sigma_{ij}+\lambda u_{i}u_{j}$, where $\lambda$ scales the advective term. Substituting the expansions from Eqs. (25) and (27) into this ansatz for the $xx$-component yields:
@@
Consequently, our effective pressure $P_{\text{eff}}\propto(1-R^{4})$ decays much more slowly with ordering than the standard $(1-R^{2})$ prediction; implying a much stiffer equation of state near ordered regions ($R\approx 1$) than standard active fluids. In our model, this creates a thermodynamic barrier: the isotropic ’disorder pressure’ remains dominant even as the system attempts to synchronize. This explains the stability of the active vortex glass (Phase II): the intermediate agitation ($\Delta\omega$) is insufficient to overcome this pressure plateau, causing the system to jam in a disordered state ($R\to 0$). Only the high agitation of Phase III provides the energy density required to penetrate this barrier and sustain the regime of marginal synchronization ($0<R<1$) characteristic of the Onsager condensate.
@@
where we, for the time being, leave the exact nature of $\mathbf{F}_{\text{active}}$ ambiguous, as we must first establish the specific hydrodynamic analogy.
VII.3 The shallow water analogy
As stated in the Methodology section, the interpretation of Eq. (18) identifies the system being modeled as isomorphic to the shallow water framework, where the active density $\rho$ acts as the fluid height.
@@ VII.3 The shallow water analogy
However, before a rigorous derivation and scrutiny of the analogy, we must point out a nuance introduced by a critical thermodynamic distinction in the equation of state. In classical shallow water systems, hydrostatic pressure scales with height ($P\propto\rho^{2}$), leading to dispersion [65]. In Eq. (31), we derived the pressure term $P_{\text{eff}}\propto(1-R^{4})$, meaning that the effective pressure scales with disorder. Crucially, because the local order parameter $R$ arises from the mean-field convolution of the particle density (Eq. 1), regions of high density support high synchronisation ($R\sim\rho$; realized by $R$ being the magnitude of the complex mean field, in turn a measurement of $\rho$, since phasors align in local synchronisation regions). Consequently, the pressure gradient $\nabla P_{\text{eff}}\propto-\nabla R^{4}$ acts as a suction force, drawing agents from disordered regions ($R\approx 0$) toward ordered vortex cores ($R\approx 1$). This manifests as a fluid with effective negative compressibility, driving the system to phase-separate into the coherent topological structures observed in the Onsager dipole, rather than dispersing into noise.
@@
Consequently, the inviscid Euler behavior that we observe in the RFE field represents the low-Mach-number limit of this compressible system, where the rotational, non-compressive wave modes dominate the irrotational, compressive wave (acoustic) modes at large scales. Hence, in the hydrodynamic limit (the Onsager dipole, or Phase III in Table 1), we recover in Eq. (32) the effective Euler advection term $\lambda(\mathbf{u}\cdot\nabla)\mathbf{u}$. Here we note that the vanishing inertia ($\lambda\to 0$) at defect cores allow for instantaneous acceleration of defects in response to potential gradients, which mimics an infinite sound speed ($c_{s}\to\infty$), physically justifying the rigid lid assumption by allowing the flow field to adjust without acoustic delay.
@@
It is instructive at this point to contrast the foregoing with the defect-driven accumulation observed in active nematics [68] and flow-stabilized vortex lattices [69]. In those systems, mass transport is driven by the divergence of an active stress that scales linearly with the nematic tensor order parameter ($\sigma\sim\mathbf{Q}$), which creates effective pressure wells that scale akin to $\propto(1-R^{2})$. In our model, however, the emergent inertia of the topological gas ($\lambda=R^{2}$) yields a stiffer effective pressure $P_{\text{eff}}\propto(1-R^{4})$ (Eq. 31). This creates sharper, ’hard-core’ density cavities at defect sites (see Figure 7 below), providing additional justification for the point vortex (Kirchhoff-Onsager) Hamiltonian in topological gas dynamics, since it assumes distinct, non-overlapping points [22].
VII.4 The Active Enthalpy Functional
Figure 7: The end-state of a Phase III (Onsager dipole) simulation, showing RFE particle locations, with colorscale denoting intrinsic frequency or frustration $\omega_{i}$ (panels a, b), the degree of local synchronization $R$ (panels c, d), and the topological charge (panels e, f). The right column shows a blown-up portion of the left column.
In this section, we shall rigorously close the derivation of the shallow water analogy, using the active enthalpy functional as a theoretical vehichle [70, 71]. However, we must first bolster the connection with a specific empirical observation. Figure 7 shows the end-state of a Onsager dipole simulation, color-coding $\omega_{i}$ (top), $R$ (middle), and topological charge $m$ (bottom). We make two important observations;
(1) As alluded to in the Main Paper, $\omega_{i}$ in Figure 7a, b) is effectively segregated into distinct bands and regions. Because of the segregation, the local average $\langle\omega_{i}\rangle$ at a position $\mathbf{r}$ becomes non-random and stable over timescales relevant to the flow. We can effectively “smear” the discrete particles into a continuous background field where $\omega(\mathbf{r})\approx\langle\omega_{i}\rangle_{\mathbf{r}}$.
(2) Figure 7a–d) present clear, unambiguous empirical evidence for the fact that, at topological defect cores, $R\to 0$ and $\omega_{i}\gg\bar{\omega}$, validating Eq. (15) and other instances where the mean field $\Psi$ appears in the numerator. As we shall motivate below, the conditions at defect cores implies the fluid is stiff. The flow must immediately adjust to the topography defined by $\omega(\mathbf{r})$. The segregation of the active particles constitutes the a potential $\Phi(\mathbf{r})$ that forces the flow.
Having motivated a conceptual shift from a Lagrangian to an Eulerian perspective, we define the effective fluid density as $\rho(\mathbf{r},t)\equiv R(\mathbf{r},t)^{2}$. We then use $\omega(\mathbf{r})$ as an external scalar potential produced by intrinsic frequency mismatch $\omega_{i}-\bar{\omega}$, $\bar{\omega}$ being mean frequency, the ensemble average of which equals dispersion $\Delta\omega$.
The total potential energy density $\mathcal{U}_{\text{tot}}$ is composed of the internal interaction energy (barotropic fluid self-interaction, or non-linear coupling in turbulence physics), and the external topographic energy [72]. We have,
$\mathcal{U}_{\text{tot}}[\rho]=\underbrace{g_{\text{eff}}/2(\rho-\rho_{0})^{2}}_{\text{Interaction}}+\underbrace{\gamma\rho\cdot\omega(\mathbf{r})}_{\text{Topography}},$ (33)
where $g_{\text{eff}}$ is the effective coupling constant representing the stiffness of the order parameter (analogous to gravity $g$ in shallow waters), and $\gamma$ is a coupling coefficient with dimensions $\text{length}^{2}\text{time}^{-1}$ relating frequency disorder $\omega(\mathbf{r})$ to potential energy. Two concise observations can now be had by dimensional analysis. (1) the units of $\gamma$ naturally lead to the definition $\gamma=\sigma v_{0}$, anchoring the topographic potential to the physical interaction scale of the agents. (2) by direct comparison between Eqs. (31) and (33), we observe that $g_{\text{eff}}\equiv v_{0}^{2}/2$.
Next, the force per unit mass (acceleration) acting on a fluid element is the negative gradient of the variational derivative of the energy with respect to density [72]:
$\mathbf{F}_{\text{active}}=-\nabla\left(\frac{\delta\mathcal{U}_{\text{tot}}}{\delta\rho}\right).$ (34)
Substituting Eq. (33) with ,
$\frac{\delta\mathcal{U}_{\text{tot}}}{\delta\rho}=\frac{v_{0}^{2}}{2}(\rho-\rho_{0})+\sigma v_{0}\omega(\mathbf{r}).$ (35)
Thus, the force field becomes,
$\mathbf{F}_{\text{active}}=-\underbrace{v_{0}^{2}\nabla\rho/2}_{\text{Barotropic Pressure}}-\underbrace{\sigma v_{0}\nabla\omega(\mathbf{r})}_{\text{Disorder Gradient}},$ (36)
where we note that $\nabla\omega$ appears as a conservative force field, akin to topographic steering in shallow waters [72], and in adaptations to active nematics [64]. Next, we will insert this force into the corrected Eulerian momentum equation for $\mathbf{u}$,
$\frac{\partial\mathbf{u}}{\partial t}+(\mathbf{u}\cdot\nabla)\mathbf{u}=\mathbf{F}_{\text{active}}+\mathbf{F}_{\text{visc}},$ (37)
where we include a viscous force term $\mathbf{F}_{\text{visc}}$ for completeness, to be addressed later. Substituting $\rho=R^{2}$ and our derived force, we obtain,
$\frac{D\mathbf{u}}{Dt}=-\frac{v_{0}^{2}}{2}\nabla(R^{2})-\sigma v_{0}\nabla\omega(\mathbf{r})+\nu\nabla^{2}\mathbf{u},$ (38)
where $\nu\nabla^{2}\mathbf{u}$ is the viscous term, and where we identified the operator $D/Dt$ in Eq. (37), the material, or total, derivative [73]. This equation now mathematically represents a fluid flowing over a variable topography,
$Z(\mathbf{r})=\frac{2\sigma}{v_{0}}\omega(\mathbf{r}),$ (39)
succinctly on display in Figure 7. Since the inertial flow necessarily moves the distribution of $\omega_{i}$, the system is effectively advecting its own boundary conditions; the active particles produce the seafloor which provides topographic steering to the emergent inertial flow.
VII.5 Active acoustic speed
At this point, we must address acoustics, as a shallow water vortex diffuses into the medium through sound waves. To demonstrate why our Onsager dipole is unable to do so, we first revisit the continuity and momentum equations (Eqs. 17 and 37) using $\rho=R^{2}$,
$\frac{\partial(R^{2})}{\partial t}+\nabla\cdot(R^{2}\mathbf{u})=0,$ (40)
and
$\frac{\partial\mathbf{u}}{\partial t}+(\mathbf{u}\cdot\nabla)\mathbf{u}=-\frac{v_{0}^{2}}{2}\nabla(R^{2})-v_{0}\sigma\nabla\omega(\mathbf{r}).$ (41)
To find the speed of acoustic waves, we linearize the system around a base state of constant synchronization $R_{0}(\mathbf{r})$ and zero velocity, neglecting the external topography $\omega(\mathbf{r})$ to isolate the acoustic modes,
$R(\mathbf{r},t)=R_{0}(\mathbf{r})+\delta R(\mathbf{r},t)$ (42)
and $\mathbf{u}=\mathbf{v}$ ($\mathbf{v}$ being a small perturbation). The linearized continuity and momentum equations then read,
$\frac{\partial\delta R(\mathbf{r},t)}{\partial t}+\frac{R_{0}(\mathbf{r})}{2}\nabla\cdot v=0,$ (43)
and
$\frac{\partial v}{\partial t}=-v_{0}^{2}R_{0}(\mathbf{r})\nabla\delta R(\mathbf{r},t),$ (44)
respectively. Combining these by taking the time derivative of the continuity equation and substituting the momentum equation yields the wave equation:
$\frac{\partial^{2}\delta R(\mathbf{r},t)}{\partial t^{2}}-\frac{v_{0}^{2}R_{0}(\mathbf{r})^{2}}{2}\nabla^{2}\delta R(\mathbf{r},t)=0,$ (45)
from which we identify the active acoustic speed $c_{s}$,
$c_{s}(R)=\frac{v_{0}R_{0}(\mathbf{r})}{\sqrt{2}}.$ (46)
In other words, the sound-speed, or, in our case, propagation speed of oscillator state information, depends linearly on the local synchronisation $R$. When an acoustic wave travels from an ordered region ($R\approx 1$) into a disordered region ($R\approx 0$), a shock front necessarily forms, since the sound-speed rapidly decreases. The Mach number ($M$) of the flow, which is defined as the ratio of flow speed to sound speed, becomes,
$M=\frac{|\mathbf{u}|}{c_{s}}\approx\frac{v_{0}R}{v_{0}R_{0}/\sqrt{2}}=\sqrt{2}\left(1+\frac{\delta R}{R_{0}}\right),$ (47)
by merit of Eqs. (14) and (46). The macroscopic flow is therefore inherently supersonic ($M>1$). Near a defect, however, $R\to 0$, and so $c_{s}\to 0$; state information cannot escape a defect core, turning the latter into sonic black holes. This implies the existence of a “sonic horizon” around defect cores, which become isolated behind Mach cones. Hence, defect cores cannot diffuse into their surroundings by radiating sound (such as geophysical vortices), justifying our appliation of the the point-vortex Onsager-Kirchoff Hamiltonian.
Figure 8: Panel a) shows acoustic speed (Eq. 46) with a colormap, with $\mathbf{u}$ overlaid with flow lines, for an end-state simulation space. Panel b) shows a synthetic Schlieren plot ($|\nabla\rho|=|\nabla R^{2}|$), showing shock ridges.
Figure 8 demonstrates the foregoing in empirical terms, showing the interplay between flow (panel a) and shock ridges (panel b) for the end-state of a Phase III (Onsager dipole) simulation. We observe that (1) the acoustic speed (colorscale in panel a) goes to zero near the defect cores, capturing the sonic black holes that are protected by Mach cones, and (2) the synthetic schlieren plot [74] indicates the presence of shocks that radiate outward from the defect cores, where they thermalize inertial mass $R^{2}$.
We are now in a position to elucidate the dynamic behaviour observed in the hydrodynamic limit of our model. The vortex merging that we observe in Figures 2 and 6, evidenced by the $t^{-0.75}$ scaling of the defect count, is explained by shock merging in the hydrodynamic limit. That is, when two synchronisation shocks collide, $R$ decreases; the inertial energy contained in $R^{2}$ must be dissipated at the shock, going back into the active bath at the long-range strain interface between the defects. The active bath again synchronises, increasing $R$ and thereby creating inertial advection, creating the remarkably stable dynamics of the Onsager dipole.
To bolster the foregoing description, we must go back to Eq. (42), we note that $\delta R(\mathbf{r},t)$ is a fast, zero-mean fluctuation $\langle\delta R\rangle_{t}=0$. We then linearise the density field $R^{2}$:
$R(\mathbf{r},t)^{2}=R_{0}(\mathbf{r})^{2}+2R_{0}(\mathbf{r})\delta R(\mathbf{r},t)+(\delta R(\mathbf{r},t))^{2}.$ (48)
When we apply the time-average operator $\langle\cdot\rangle_{t}$ to this equation, the linear fluctuation term vanishes, and we are left with $\langle R^{2}\rangle_{t}\approx R_{0}(\mathbf{r})^{2}$.
Next, we define the active velocity potential (or macroscopic phase) $\Theta(\mathbf{r},t)$ [63],
$\Theta(\mathbf{r},t)=\bar{\Theta}(\mathbf{r},t)+\delta\Theta(\mathbf{r},t),$ (49)
where a Helmholtz decomposition yields slow topological rotation (vortices) and the fast compressional vibration (acoustic waves). Our goal is to show that the slow-mode is incompressible, yielding the stable Euler dynamics that we observe in the main paper. $\delta\Theta$ describes the irrotational phase oscillations associated with the sound waves $\delta R$, and likewise time-average to zero: $\langle\delta\Theta\rangle_{\tau}=0$. We substitute both expansions into the continuity equation,
$\partial_{t}R(\mathbf{r},t)^{2}+\nabla\cdot(R(\mathbf{r},t)^{2}\nabla\Theta(\mathbf{r},t))=0,$ (50)
and time-average,
$\nabla\cdot\langle(R_{0}(\mathbf{r})^{2}+2R_{0}(\mathbf{r})\delta R(\mathbf{r},t))\nabla(\bar{\Theta}(\mathbf{r},t)+\delta\Theta(\mathbf{r},t))\rangle_{\tau}\approx 0.$ (51)
After expanding Eq. (51), the second and third terms vanish by merit of the time-avering, while the fourth term vanish as $\delta R(\mathbf{r},t)$ and $\nabla\delta\Theta(\mathbf{r},t)$ are uncorrelated. We are then left with the first dominant term, and recover the elliptic constraint for the vortex phase $\bar{\Theta}$ on the background $R_{0}$,
$\nabla\cdot(R_{0}^{2}\nabla\bar{\Theta})=0,$ (52)
where we drop spatial and temporal dependencies for brevity. Eq. (52) describes an incompressible potential flow in a medium with variable density $R_{0}^{2}$, and is analogous to the rigid lid approximation [63]. We recognise a stable Euler fluid with inertial mass $R_{0}^{2}$, flowing around defect cores, and Eq. (52 enforce mass conservation for the inertial background flow.
To address the physical nature of the background flow (which, at present, remains a static field approximation), we note that while the external topography $\omega(\mathbf{r})$ was neglected to isolate the fast acoustic modes ($\delta R$), it remains the governing factor for the slow background state $R_{0}(\mathbf{r})$. The static balance of the active enthalpy functional (Eq. 33) dictates that $R_{0}(\mathbf{r})^{2}$ is spatially anti-correlated with $\omega(\mathbf{r})$, such that high-disorder regions create the density depressions (vacuums) that isolate defects, while low-disorder regions support the high-density ($R_{0}\approx 1$) inertial fluid. A close examination of Figure 7b, d) provides clear empirical evidence for this position.
VII.6 Hydrodynamic closure
Finally, we reconcile the stiffness of the effective pressure ($P_{\text{eff}}\propto 1-R^{4}$) with the observation of compressible shocks. As the active acoustic speed (Eq. 46) scales linearly with order ($c_{s}\propto R$), the stiffness ensures that the sound speed inherently lags behind the inertial flow velocity ($|\mathbf{u}|\propto R$, Eq. 14), forcing the system into a global supersonic state with $M\approx\sqrt{2}$. As we discussed in the foregoing subsection, this condition mandates the formation of synchronization shocks (Mach cones) at defect sites, which act as the primary mechanism topological sorting (yielding Eulerian Onsager-Kirchoff dynamics in the supersonic topological gas). This leads us to assert the hazardous rigid lid incompressibility,
$\nabla\cdot\mathbf{u}\approx 0.$ (53)
The “phantom inertia” $R^{2}$ acts an Euler fluid with quantized circulation, which we demonstrate with rigour in the Main Paper. At the same time, on fast, acoustic timescales, the dynamics are dominated by compressible shock-mergers. The shocks thermalise inertial energy into the active bath, which subsequently build up order through inverse cascading enthropy. The rigid lid incompressibility emerges as the rigorous time-averaged limit of the stiff, supersonic topological gas.